OpenStudy (anonymous):

road trip total 48 1/3 miles. 25 4/9 already traveld, how many miles left? how do i calculate S.O.S

OpenStudy (anonymous):

answer is on the bottom. hope this is what you need http://mathway.com/answer.aspx?p=calg?p=SMB15(48(1SMB103))-(25(4SMB109))?p=2?p=?p=?p=?p=?p=0?p=?p=0?p=?p=

OpenStudy (gw2011):

Find a common denominator which in this case is 9. 48 1/3=48 3/9=47 12/9 The reason this is 47 12/9 is because 25 4/9 will be subtracted and 4/9 is larger than 3/9. Therefore, you borrow 1 or 9/9 from 48 and add it to 3/9 to get 12/9. 47 12/9 - 25 4/9 = 22 8/9

OpenStudy (anonymous):

thaaaaank u again =))

OpenStudy (gw2011):

You're welcome